# Bayesian Inference of Gravity through Realistic 3D Modeling of Wide Binary Orbits: General Algorithm and a Pilot Study with HARPS Radial Velocities

### v4

@@ Abstract
When 3D relative displacement $\mathbf{r}$ and velocity $\mathbf{v}$ between the pair in a gravitationally-bound system are precisely measured, the six measured quantities at one phase can allow elliptical orbit solutions at a given gravitational parameter $G$. Due to degeneracies between orbital-geometric parameters and $G$, individual Bayesian inferences and their statistical consolidation are needed to infer $G$ as recently suggested by a Bayesian 3D modeling algorithm. Here I present a fully general Bayesian algorithm suitable for wide binaries with two (almost) exact sky-projected relative positions (as in the Gaia data release 3) and the other four sufficiently precise quantities. Wide binaries meeting the requirements of the general algorithm to allow for its full potential are rare at present, largely because the measurement uncertainty of the line-of-sight (radial) separation is usually larger than the true separation. As a pilot study, the algorithm is applied to 32 Gaia binaries for which precise HARPS radial velocities are available. The value of $\Gamma\equiv\log_{10}\sqrt{G/G_{\rm N}}$ (where $G_{\rm N}$ is Newton’s constant) is $0.002_{-0.018}^{+0.012}$ supporting Newton for a combination of 24 binaries with Newtonian acceleration $g_{\rm N}>10^{-9}$m s-2, while it is $\Gamma=0.134_{-0.036}^{+0.056}$ (thermal prior on eccentricity) or $0.115_{-0.028}^{+0.048}$ (flat prior) for 8 binaries with $g_{\rm N}<10^{-9}$m s-2 representing $3.7\sigma$ discrepancy with Newton. Moreover, the inferred orbital true anomalies clearly favor modified gravity over Newton with the difference of Bayesian information criterion >10>10. The pilot study demonstrates the potential of the algorithm in measuring and testing gravity at low acceleration with future samples of wide binaries.
@@ 1 Introduction
Gravitationally-bound binary systems have played pivotal roles in astronomy and physics. The orbital motion of planets in the Sun-planet binary systems provided the early experimental foundation of Newtonian gravitational dynamics. The gravitational anomaly in Mercury’s orbital motion in a relatively strong gravity regime provided an early experimental success of Einstein’s general relativity. Binary (and higher-multiplicity) stars have been playing fundamental roles in determining stellar masses and radii and constraining star formation and evolution models (e.g., Duchene:2013; ElBadry:2024).
@@
A currently interesting use of binary stars in studies of fundamental physics is to probe gravity in the low-acceleration ($\lesssim 10^{-9}$m s-2) regime with wide binaries. The dark matter (DM) phenomena in galaxies indicate a critical acceleration scale $a_{0}\approx 1.2\times 10^{-10}$m s-2, known as Milgrom’s or modified Newtonian Dynamics (MOND; Milgrom:1983) constant (see the reviews by Sanders:2002; Famaey:2012). As Milgrom proposed, the presence of $a_{0}$ in galaxies may indicate the breakdown of standard Newton-Einstein gravity in the low-acceleration regime, opening possibilities of gravity theories breaking the strong equivalence principle (Milgrom:1983). This can be directly tested by wide binaries in the solar neighborhood as wide binary internal dynamics cannot be affected by the dark matter density inferred for the Milky Way (Hernandez:2012).
@@
There is a fundamental limitation in probing gravity with individual wide binaries because those of interest have long periods of $\gtrsim 10^{5}$ yr and thus their orbits cannot be traced for significant segments. When the relative displacement $\mathbf{r}$ and the relative velocity $\mathbf{v}$ between the pair are measured with sufficient precision, the six components at one phase can be used to reconstruct the orbit assuming Newtonian gravitational dynamics (see, e.g., Kervella:2017 and Akeson:2021 for reconstructed orbits of Proxima around $\alpha$ Cen AB). However, the same data may be equally well fitted by nonstandard theories of gravity with varied orbits. In other words, globally different orbits for different theories of gravity may fit the instantaneous displacement and velocity equally well. Thus, such degeneracies must be dealt with in using wide binaries for testing gravity through 3D orbit modeling. In the specific case of Proxima, its exceptional proximity to the Sun may allow a future monitoring campaign to break the degeneracy between standard and nonstandard gravities (BanikZhao:2018; BanikKroupa:2019).
@@
Wide binaries with precise $\mathbf{r}$ and $\mathbf{v}$ are rare because it is challenging to measure the relative radial (line-of-sight) displacement and the relative radial velocity (RV) with sufficiently good precision. In particular, the measurement uncertainties of the relative distances are usually (much) larger than the radial separations, whereas precise RVs are currently being obtained for increasingly larger samples. Nevertheless, the potential importance of 3D velocities $\mathbf{v}$ with precise RVs was recognized by earlier investigators of wide binaries (Scarpa:2017; PittordisSutherland:2018).
@@
Recently, two studies considered 3D orbit modeling to test gravity based on samples of limited quantity and/or quality. For 32 binaries selected from the Gaia data release 3 (DR3; Gaia:2023) database, Saglia:2025 derived very precise RVs from the High Accuracy Radial Velocity Planet Searcher (HARPS) spectrograph database. Their sample covers a broad range of Newtonian gravitational acceleration $10^{-11}\lesssim g_{\rm N}\lesssim 10^{-6}$ m s-2 including about 8 wide binaries with $g_{\rm N}<10^{-9}$ m s-2 as will be shown below. Notably, their sample includes only a few binaries with uncertainty of distance difference ($\Delta d$) satisfying $\sigma_{\Delta d}<s/\sqrt{2}$ where $s$ is the sky-projected 2D separation. Thus, individual orbit inferences and gravity tests are limited by overall imprecise radial separations (note that tangential separations derived from Gaia astrometry have negligibly small errors). Moreover, even if $\sigma_{\Delta d}$ is sufficiently small, degeneracies between orbits and gravity theories need to be dealt with. Saglia:2025 presented only non-unique Newtonian orbit solutions for all their binaries except for one system (#24) with stars HD189739 and HD189760 which they interpret as an unbound system. They have not derived uncertainties or probability distributions of orbit parameters, but they noted that some orbit phases seemed unnatural.
@@
Chae:2025 introduced a Bayesian methodology to infer for each binary a posterior probability distribution of (nuisance) orbit parameters along with the gravitational boost (or anomaly) parameter
@@
where $G$ is the generalized gravitational parameter and $G_{\rm N}$ is Newton’s constant. The individual PDFs of $\Gamma$ are then statistically consolidated (Hill:2011) for binaries having similar internal gravitational accelerations to derive $\Gamma$ as a function of acceleration. The methodology was applied to a sample of 312 wide binaries with relative RV uncertainties <380<380m s-2 selected from the Gaia DR3. One caveat of Chae:2025 was that the orbit configuration suppressed one of the two parameters representing the direction of the orbit axis with respect to the observer’s view, in part because the observational constraints were not sufficient. For this reason, the Chae:2025 methodology will be referred to as ‘simple Bayesian’ throughout.
The purpose of this work is two-fold. One is to present a fully general version of the Chae:2025 methodology that can be used for wide binaries with sufficiently precise measurements of all components of $\mathbf{r}$ and $\mathbf{v}$. At present, such wide binaries are extremely rare, but it is in principle more correct to use a fully general model even for insufficient data (although the parameters may be less well constrained). The other is to to apply the methodology to the Saglia:2025 sample with extremely precise HARPS RVs, as a pilot study expecting (much) larger samples of similar qualities in the future. This pilot study can also complement Saglia:2025 results by deriving Bayesian probability distributions of not only orbit and orientation parameters but also $\Gamma$. However, note that except for a few binaries of the Saglia:2025 sample, $\Gamma$ and the orbit parameters are not well constrained individually. The Python codes implementing the general Bayesian algorithm (Chae:Zenodo2025) and the Bayesian outputs for the Saglia:2025 sample are available on Zenodo under an open-source Creative Commons Attribution license: 10.5281/zenodo.17113129.
2 Methodology
Following Chae:2025 I work only with elliptical orbits, meaning that Newtonian or pseudo-Newtonian gravity with a generalized gravitational parameter is considered. In this approach, any nonstandard gravity is approximated to be pseudo-Newtonian in wide binary internal dynamics.
@@ 2 Methodology
For a binary of stars $A$ (brighter) and $B$ with an observational total mass $M_{\rm tot}\equiv M_{A}+M_{B}$ and sky-projected separation $s\equiv\sqrt{(x^{\prime})^{2}+(y^{\prime})^{2}}$, we have the magnitude of the 3D separation
@@
where $f_{M}$ is a factor to take into account the uncertainty of the observational mass. Note that parameter $a$ is not explicitly used in Equations (2) and (3).
@@
from which follows $\phi=\tan^{-1}(b/\cos i)$ (Equation (14) of Chae:2025) when $\tan\theta=0$. Then, we have the following prediction on the four remaining quantities $z^{\prime}$, $v_{x^{\prime}}$, $v_{y^{\prime}}$, and $v_{z^{\prime}}$ expressed as functions of $e$, $i$, $\phi_{0}$, and $\phi$ only:
@@
where $k(=\pm 1)$ indicates the sign of the quantity $dx/d\phi$ or ($\sin\phi+e\sin\phi_{0}$). Equations (6) and (7) follow from Equations (A5) and (A6) of Chae:2025.
@@
Priors on the model parameters are considered depending on how gravity is probed/tested with the Bayesian methodology. Two approaches are considered. In the first approach, as in Saglia:2025, gravity is fixed for each wide binary and the model parameters are constrained with or without priors, and the distributions of the inferred parameters are tested against the expectations for the assumed gravity. This approach will be considered in this work not only for Newtonian gravity, but also for pseudo-Newtonian gravities (where $G$ is allowed to vary from system to system depending on the internal Newtonian acceleration) inspired by existing MOND gravity models. In the second approach, as in Chae:2025, logical or empirical priors are imposed on all the free parameters except for $\Gamma$, and PDFs of $\Gamma$ are derived. For the purpose of inferring/testing gravity, this will be the main approach of this work.
@@
Further priors are considered for mass parameter $f_{M}$ and eccentricity $e$. Because the total mass of each system was determined observationally, I consider a lognormal probability distribution of $f_{M}$ with mean =0=0 and s.d. $5\%$. This prior on $f_{M}$ is always imposed regardless of the other priors. Because a very precise 3D velocity has some constraining power on $e$, a prior on $e$ is not a significant factor (unlike Chae:2025). I consider two cases: a generic distribution $f_{\text{pr}}(e)=2e$ (the thermal probability distribution) and the flat/no prior within the range $0<e<1$. Finally, when $\Gamma$ is constrained for each binary, it is assumed to be in the range $1<\Gamma<1$, which turns out to include most PDFs sufficiently well.
@@
Markov chain Monte Carlo (MCMC) procedure is performed to derive the posterior distribution $p(\mathbf{\Theta})$ with the public Python package emcee (emcee). The numerical procedure is as follows. First, approximate global best-fit values of $e$, $i$, $\phi_{0}$, $\Delta\phi$, and $\Gamma$ are determined by a grid search of the parameter space to minimize $\chi^{2}=\Sigma_{j}((X_{j,\rm{mod}}(\mathbf{\Theta})-X_{j,\rm{obs}})/{\sigma_{j}})^{2}$ (see Equation (12) for the definition of symbols). These values along with $\log_{10}f_{M}=0$ are the starting values of random walks. The number of random walkers is nwalkers=200=200. With six parameters (ndim=6=6) that is one more than the simple Bayesian approach of Chae:2025, I consider longer chains of iterations than Chae:2025. The particular choice to derive the nominal results is the following: niter=400000=400000, discard=200000=200000 (to well exclude the burn-in phase), and thin=20=20, so that a total of nwalkers $\times$ (niter-discard) // thin $2\times 10^{6}$ models are obtained to derive posterior probability density functions (PDFs) of the parameters.
@@
3 A pilot study with HARPS radial velocities: Results and Discussion
@@
For each binary from the Saglia:2025 sample, Bayesian results are produced based on the two approaches described in Section 2: (1) fixed gravity with or without priors on the other parameters (except for the mass parameter $f_{M}$ whose prior is always imposed); (2) free $\Gamma$ with priors on the other parameters. In the first approach, two cases are considered. In one case, gravity is Newtonian for all 32 binaries. In the other case, hereafter to be referred to as the “MOND-type toy model”, $G$ is fixed for each binary at a value depending on its internal Newtonian acceleration (inferred from the Newtonian model) $g_{\rm N}$ as follows: $G=G_{\rm N}$ for $g_{\rm N}>10^{-9}$ m s-2 and $G=1.4G_{\rm N}$ for $g_{\rm N}<10^{-9}$ m s-2. This toy model is clearly not a whole theoretical prediction of MOND or any other modified gravity. It is introduced here to serve as a control nonstandard model (or a working approximation) that is intended to mimic a transition from the verified Newtonian regime of $g_{\rm N}\gtrsim 10^{-8}$ m s-2 (Chae:2025) to the low-acceleration regime of $g_{\rm N}<10^{-9}$ m s-2. Moreover, as will be shown in this work, a subsample of 23 binaries with $g_{\rm N}>10^{-9}$ m s-2 agrees well with Newton while the other subsample does not. Thus, this toy model is self-consistent within this work.
@@
The relative RV between the two stars $v_{z^{\prime}}=-(\text{RV}_{B}-\text{RV}_{A})$ used here is obtained by correcting RVs taken from Table B.2 of Saglia:2025 for gravitational redshift and convective flow. The velocity components are corrected for perspective effects (Shaya:2011; Yoon:2025) because some binaries are close to the Sun (although they are in general negligibly small). The nominal uncertainty of $v_{z^{\prime}}$ is enlarged by quadratically adding $40$ m s-1 to it following the recommendation of Saglia:2025.
@@
While $x^{\prime}_{\rm obs}$ and $y^{\prime}_{\rm obs}$ are measured with negligibly small errors, $z^{\prime}_{\rm obs}=-(d_{B}-d_{A})$ is poorly measured in general. Thus, even if all the velocity components are very precise, we do not expect that the model parameters are well constrained, especially for the generalized gravity model. Considering that the expected radial separation is of the order of $s/\sqrt{2}$, the required precision is $\sigma_{z^{\prime}}<s/\sqrt{2}$ for $S/N>1$. Only three systems meet this requirement: #5, #12, and #21 from Table B.1 of Saglia:2025. However, note that even for these systems, the precision of $z^{\prime}_{\rm obs}$ may not be sufficiently good. Thus, unfortunately, in general we do not expect narrow posterior ranges of the parameters.
@@
3.1 Distribution of the Bayesian Inferred Parameters

### v5

@@ Abstract
When 3D relative displacement $\mathbf{r}$ and velocity $\mathbf{v}$ between the pair in a gravitationally-bound system are precisely measured, the six measured quantities at one phase can allow elliptical orbit solutions at a given gravitational parameter $G$. Due to degeneracies between orbital-geometric parameters and $G$, individual Bayesian inferences and their statistical consolidation are needed to infer $G$ as recently suggested by a Bayesian 3D modeling algorithm. Here I present a fully general Bayesian algorithm suitable for wide binaries with two (almost) exact sky-projected relative positions (as in the Gaia data release 3) and the other four sufficiently precise quantities. Wide binaries meeting the requirements of the general algorithm to allow for its full potential are rare at present, largely because the measurement uncertainty of the line-of-sight (radial) separation is usually larger than the true separation. As a pilot study, the algorithm is applied to 32 Gaia binaries for which precise HARPS radial velocities are available. The value of $\Gamma\equiv\log_{10}\sqrt{G/G_{\rm N}}$ (where $G_{\rm N}$ is Newton’s constant) is $0.002_{-0.018}^{+0.012}$ supporting Newton for a combination of 24 binaries with Newtonian acceleration $g_{\rm N}>10^{-9}$m s-2, while it is $\Gamma=0.134_{-0.036}^{+0.056}$ ($0.143_{-0.041}^{+0.068}$) for 8 (6) binaries with $g_{\rm N}<10^{-9}$ ($10^{-9.5}$) m s-2 representing $\gtrsim 3.5\sigma$ discrepancy with Newton. However, one system (Stars HD189739 and HD189760) dominates the signal. Without it, the tension with Newton is significantly lessened with $\Gamma=0.063_{-0.041}^{+0.065}$. Thus, to verify the tentative signal, many such systems need to be discovered and their kinematic nature such as any possibility of hidden tertiary stars needs to be thoroughly addressed. The pilot study demonstrates the potential of the algorithm in measuring and testing gravity at low acceleration with future samples of wide binaries.
@@ I Introduction
Gravitationally-bound binary systems have played pivotal roles in astronomy and physics. The orbital motion of planets in the Sun-planet binary systems provided the early experimental foundation of Newtonian gravitational dynamics. The gravitational anomaly in Mercury’s orbital motion in a relatively strong gravity regime provided an early experimental success of Einstein’s general relativity. Binary (and higher-multiplicity) stars have been playing fundamental roles in determining stellar masses and radii and constraining star formation and evolution models (e.g., Duchêne and Kraus 2013; El-Badry 2024).
@@
A currently interesting use of binary stars in studies of fundamental physics is to probe gravity in the low-acceleration ($\lesssim 10^{-9}$m s-2) regime with wide binaries. The dark matter (DM) or missing gravity phenomena in galaxies indicate a critical acceleration scale $a_{0}\approx 1.2\times 10^{-10}$m s-2, known as Milgrom’s or modified Newtonian Dynamics (MOND; Milgrom, 1983) constant (see the reviews by Sanders and McGaugh 2002; Famaey and McGaugh 2012; Banik and Zhao 2022). As Milgrom proposed, the presence of $a_{0}$ in galaxies may indicate the breakdown of standard Newton-Einstein gravity in the low-acceleration regime, opening possibilities of gravity theories breaking the strong equivalence principle (Milgrom, 1983). This can be directly tested by wide binaries in the solar neighborhood as wide binary internal dynamics cannot be affected by the dark matter density inferred for the Milky Way (Hernandez et al. 2012; see also, e.g., Acedo 2020) while the phantom dark matter density around a small-scale star is much higher than that of the Milky Way (see, e.g., Pflamm-Altenburg 2025).
@@
There is a fundamental limitation in probing gravity with individual wide binaries because those of interest have long periods of $\gtrsim 10^{5}$ yr and thus their orbits cannot be traced for significant segments. When the relative displacement $\mathbf{r}$ and the relative velocity $\mathbf{v}$ between the pair are measured with sufficient precision, the six components at one phase can be used to reconstruct the orbit assuming Newtonian gravitational dynamics (see, e.g., Kervella et al. (2017) and Akeson et al. (2021) for reconstructed orbits of Proxima around $\alpha$ Cen AB). However, the same data may be equally well fitted by nonstandard theories of gravity with varied orbits. In other words, globally different orbits for different theories of gravity may fit the instantaneous displacement and velocity equally well. Thus, such degeneracies must be dealt with in using wide binaries for testing gravity through 3D orbit modeling. In the specific case of Proxima, its exceptional proximity to the Sun may allow a future monitoring campaign to break the degeneracy between standard and nonstandard gravities (Banik and Zhao, 2018; Banik and Kroupa, 2019).
@@
Wide binaries with precise $\mathbf{r}$ and $\mathbf{v}$ are rare because it is challenging to measure the relative radial (line-of-sight) displacement and the relative radial velocity (RV) with sufficiently good precision. In particular, the measurement uncertainties of the relative distances are usually (much) larger than the radial separations, whereas precise RVs are currently being obtained for increasingly larger samples. Nevertheless, the potential importance of 3D velocities $\mathbf{v}$ with precise RVs was recognized by earlier investigators of wide binaries (Scarpa et al., 2017; Pittordis and Sutherland, 2018).
@@
Recently, two studies considered 3D orbit modeling to test gravity based on samples of limited quantity and/or quality. For 32 binaries selected from the Gaia data release 3 (DR3; Vallenari et al., 2023) database, Saglia et al. (2025) derived very precise RVs from the High Accuracy Radial Velocity Planet Searcher (HARPS) spectrograph database. Their sample covers a broad range of Newtonian gravitational acceleration $10^{-11}\lesssim g_{\rm N}\lesssim 10^{-6}$ m s-2 including about 8 wide binaries with $g_{\rm N}<10^{-9}$ m s-2 as will be shown below. Notably, their sample includes only a few binaries with uncertainty of distance difference ($\Delta d$) satisfying $\sigma_{\Delta d}<s/\sqrt{2}$ where $s$ is the sky-projected 2D separation. Thus, individual orbit inferences and gravity tests are limited by overall imprecise radial separations (note that tangential separations derived from Gaia astrometry have negligibly small errors). Moreover, even if $\sigma_{\Delta d}$ is sufficiently small, degeneracies between orbits and gravity theories need to be dealt with. Saglia et al. (2025) presented only non-unique Newtonian orbit solutions for all their binaries except for one system (#24) with stars HD189739 and HD189760 which they interpret as an unbound system. They have not derived uncertainties or probability distributions of orbit parameters, but they noted that some orbit phases seemed unnatural.
@@
Chae (2025b) introduced a Bayesian methodology to infer for each binary a posterior probability distribution of (nuisance) orbit parameters along with the gravitational boost (or anomaly) parameter
@@
where $G$ is the generalized gravitational parameter and $G_{\rm N}$ is Newton’s constant. The individual PDFs of $\Gamma$ are then statistically consolidated (Hill and Miller, 2011) for binaries having similar internal gravitational accelerations to derive $\Gamma$ as a function of acceleration. The methodology was applied to a sample of 312 wide binaries with relative RV uncertainties <380<380m s-2 selected from the Gaia DR3. One caveat of Chae (2025b) was that the orbit configuration suppressed one of the two parameters representing the direction of the orbit axis with respect to the observer’s view, in part because the observational constraints were not sufficient. For this reason, the Chae (2025b) methodology will be referred to as ‘simple Bayesian’ throughout.
The purpose of this work is two-fold. One is to present a fully general version of the Chae (2025b) methodology that can be used for wide binaries with sufficiently precise measurements of all components of $\mathbf{r}$ and $\mathbf{v}$. At present, such wide binaries are extremely rare, but it is in principle more correct to use a fully general model even for insufficient data (although the parameters may be less well constrained). The other is to to apply the methodology to the Saglia et al. (2025) sample with extremely precise HARPS RVs, as a pilot study expecting (much) larger samples of similar qualities in the future. This pilot study can also complement Saglia et al. (2025) results by deriving Bayesian probability distributions of not only orbit and orientation parameters but also $\Gamma$. However, note that except for a few binaries of the Saglia et al. (2025) sample, $\Gamma$ and the orbit parameters are not well constrained individually. The Python codes implementing the general Bayesian algorithm (Chae, 2025a) and the Bayesian outputs for the Saglia et al. (2025) sample are available on Zenodo under an open-source Creative Commons Attribution license: 10.5281/zenodo.17113129.
II Methodology
Following Chae (2025b) I work only with elliptical orbits, meaning that Newtonian or pseudo-Newtonian gravity with a generalized gravitational parameter is considered. In this approach, any nonstandard gravity is approximated to be pseudo-Newtonian in wide binary internal dynamics. This assumption is completely valid in testing Newtonian gravity and quantifying a degree of deviation from it if there arises a discrepancy with Newton. However, this assumption is approximate in testing a specific nonstandard gravity model because its predicted numerical orbits may not be closed ellipses.
@@ II Methodology
For a binary of stars $A$ (brighter) and $B$ with an observational total mass $M_{\rm tot}\equiv M_{A}+M_{B}$ and sky-projected separation $s\equiv\sqrt{(x^{\prime})^{2}+(y^{\prime})^{2}}$, we have the magnitude of the 3D displacement
@@
where $f_{M}$ is a factor to take into account the uncertainty of the observational mass. Note that variables $r$ and $a$ appearing in Equations (2) and (3) depend on the other variables in Equation (2), i.e., the directly measurable quantity $s$ and the independent variables $\phi$, $\phi_{0}$, $i$, and $e$.
@@
from which follows $\phi=\tan^{-1}(b/\cos i)$ (Equation (14) of Chae (2025b)) when $\tan\theta=0$. Then, we have the following prediction on the four remaining quantities $z^{\prime}$, $v_{x^{\prime}}$, $v_{y^{\prime}}$, and $v_{z^{\prime}}$ expressed as functions of $e$, $i$, $\phi_{0}$, and $\phi$ only:
@@
where $k(=\pm 1)$ indicates the sign of the quantity $dx/d\phi$ or ($\sin\phi+e\sin\phi_{0}$). Equations (6) and (7) follow from Equations (A5) and (A6) of Chae (2025b).
@@
Priors on the model parameters are considered depending on how gravity is probed/tested with the Bayesian methodology. Two approaches are considered. In the first approach, as in Saglia et al. (2025), gravity is fixed for each wide binary and the model parameters are constrained with or without priors, and the distributions of the inferred parameters are tested against the expectations for the assumed gravity. This approach will be considered in this work not only for Newtonian gravity, but also for pseudo-Newtonian gravities (where $G$ is allowed to vary from system to system depending on the internal Newtonian acceleration) inspired by existing MOND gravity models. In the second approach, as in Chae (2025b), logical or empirical priors are imposed on all the free parameters except for $\Gamma$, and PDFs of $\Gamma$ are derived. For the purpose of inferring/testing gravity, this will be the main approach of this work.
@@
Further priors are considered for mass parameter $f_{M}$ and eccentricity $e$. Because the total mass of each system was determined observationally, I consider a lognormal probability distribution of $f_{M}$ with mean =0=0 and s.d. $5\%$. This prior on $f_{M}$ is always imposed regardless of the other priors. Because a very precise 3D velocity has some constraining power on $e$, a prior on $e$ is not a significant factor (unlike Chae 2025b). I consider two cases: a generic distribution $f_{\text{pr}}(e)=2e$ (the thermal probability distribution) and the flat/no prior within the range $0<e<1$. Finally, when $\Gamma$ is constrained for each binary, it is assumed to be in the range $1<\Gamma<1$, which turns out to include most PDFs sufficiently well.
@@
Markov chain Monte Carlo (MCMC) procedure is performed to derive the posterior distribution $p(\mathbf{\Theta})$ with the public Python package emcee (Foreman-Mackey et al., 2013). The numerical procedure is as follows. First, approximate global best-fit values of $e$, $i$, $\phi_{0}$, $\Delta\phi$, and $\Gamma$ are determined by a grid search of the parameter space to minimize $\chi^{2}=\Sigma_{j}((X_{j,\rm{mod}}(\mathbf{\Theta})-X_{j,\rm{obs}})/{\sigma_{j}})^{2}$ (see Equation (12) for the definition of symbols). These values along with $\log_{10}f_{M}=0$ are the starting values of random walks. The number of random walkers is nwalkers=200=200. With six parameters (ndim=6=6) that is one more than the simple Bayesian approach of Chae (2025b), I consider longer chains of iterations than Chae (2025b). The particular choice to derive the nominal results is the following: niter=400000=400000, discard=200000=200000 (to well exclude the burn-in phase), and thin=20=20, so that a total of nwalkers $\times$ (niter-discard) // thin $2\times 10^{6}$ models are obtained to derive posterior probability density functions (PDFs) of the parameters.
@@
III A pilot study with HARPS radial velocities: Results and Discussion
@@
For each binary from the Saglia et al. (2025) sample, Bayesian results are produced based on the two approaches described in Section II: (1) fixed gravity with or without priors on the other parameters (except for the mass parameter $f_{M}$ whose prior is always imposed); (2) free $\Gamma$ with priors on the other parameters. In the first approach, two cases are considered. In one case, gravity is Newtonian for all 32 binaries. In the other case, hereafter to be referred to as the “MOND-type toy model”, $G$ is fixed for each binary at a value depending on its internal Newtonian acceleration (inferred from the Newtonian model) $g_{\rm N}$ as follows: $G=G_{\rm N}$ for $g_{\rm N}>10^{-9}$ m s-2 and $G=1.4G_{\rm N}$ for $g_{\rm N}<10^{-9}$ m s-2. This toy model is clearly not a whole theoretical prediction of MOND or any other modified gravity. It is introduced here to serve as a control nonstandard model (or a working approximation) that is intended to mimic a transition from the verified Newtonian regime of $g_{\rm N}\gtrsim 10^{-8}$ m s-2 (Chae, 2025b) to the low-acceleration MOND regime of $g_{\rm N}\lesssim a_{0}$. Moreover, as will be shown in this work, a subsample of 23 binaries with $g_{\rm N}>10^{-9}$ m s-2 agrees well with Newton while the other subsample does not. Thus, this toy model is self-consistent within this work.
@@
The relative RV between the two stars $v_{z^{\prime}}=-(\text{RV}_{B}-\text{RV}_{A})$ used here is obtained by correcting RVs taken from Table B.2 of Saglia et al. (2025) for gravitational redshift and convective flow. The velocity components are corrected for perspective effects (Shaya and Olling, 2011; Yoon et al., 2025) because some binaries are close to the Sun (although they are in general negligibly small). The nominal uncertainty of $v_{z^{\prime}}$ is enlarged by quadratically adding $40$ m s-1 to it following the recommendation of Saglia et al. (2025).
@@
While $x^{\prime}_{\rm obs}$ and $y^{\prime}_{\rm obs}$ are measured with negligibly small errors, $z^{\prime}_{\rm obs}=-(d_{B}-d_{A})$ is poorly measured in general. [^2] See Section 2.3 of Banik (2019) for a statistical inference of the radial separation that may be applicable for statistical studies with a large sample. Thus, even if all the velocity components are very precise, we do not expect that the model parameters are well constrained, especially for the generalized gravity model. Considering that the expected radial separation is of the order of $s/\sqrt{2}$, the required precision is $\sigma_{z^{\prime}}<s/\sqrt{2}$ for $S/N>1$. Only three systems meet this requirement: #5, #12, and #21 from Table B.1 of Saglia et al. (2025). However, note that even for these systems, the precision of $z^{\prime}_{\rm obs}$ may not be sufficiently good. Thus, unfortunately, in general we do not expect narrow posterior ranges of the parameters.
@@
III.1 Distribution of the Bayesian Inferred Parameters